# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ C.2  Oracle inequality of the Doob’s $h$-function matching
    $\leq\frac{60\lambda^{2}d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}+\frac{1}{2}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\frac{8\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t}),$
Step 1. Since ${\mathscr{H}}_{t}$ is convex, for each $h_{t}\in{\mathscr{H}}_{t}$, we have $\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda})\in{\mathscr{H}}_{t}$ for each $\varepsilon\in(0,1)$. Since that $\widehat{h}_{t}^{\lambda}$ is the minimizer of the empirical risk $\widehat{{\mathcal{J}}}_{t}^{\lambda}$ over ${\mathscr{H}}_{t}$, it follows that
$\frac{1}{\varepsilon}\Big\{\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda}))-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})\Big\}\geq 0.$
Taking the limit as $\varepsilon\rightarrow 0^{+}$ on both sides of this inequality yields
$\delta\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda},h_{t}-\widehat{h}_{t}^{\lambda})$
$\coloneq\lim_{\varepsilon\rightarrow 0^{+}}\frac{1}{\varepsilon}\Big\{\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda}))-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})\Big\}$
$=\frac{\,\mathrm{d}}{\,\mathrm{d}\varepsilon}\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda}))\Big|_{\varepsilon=0}$
$=\frac{\,\mathrm{d}}{\,\mathrm{d}\varepsilon}\widehat{{\mathcal{J}}}_{t}(\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda}))\Big|_{\varepsilon=0}+\lambda\frac{\,\mathrm{d}}{\,\mathrm{d}\varepsilon}\widehat{{\mathcal{G}}}_{t}(\widehat{h}_{t}^{\lambda}+\varepsilon(h_{t}-\widehat{h}_{t}^{\lambda}))\Big|_{\varepsilon=0}$
$=\frac{2}{n}\sum_{i=1}^{n}\Big\{(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{t}^{i})-w(\mathbf{X}_{0}^{i}))(h_{t}(\mathbf{X}_{t}^{i})-\widehat{h}_{t}^{\lambda}(\mathbf{X}_{t}^{i}))+\lambda\langle\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{t}^{i}),\nabla h_{t}(\mathbf{X}_{t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{t}^{i})\rangle\Big\}$
$\geq 0,$
where $\delta\widehat{{\mathcal{J}}}_{t}^{\lambda}$ denotes the first variation [Giaquinta and Hildebrandt, 2004] of $\widehat{{\mathcal{J}}}_{t}^{\lambda}$. Consequently,
(C.9) $\underbrace{\frac{2}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))(h_{t}(\mathbf{X}_{T-t}^{i})-\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i}))}_{\text{(i)}}$
$+\underbrace{\frac{2\lambda}{n}\sum_{i=1}^{n}\langle\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle}_{\text{(ii)}}$
$\geq-\frac{2\lambda}{n}\sum_{i=1}^{n}\nabla\langle h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle.$
For term (i) in the right-hand side of (C.9), we have
$\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))(h_{t}(\mathbf{X}_{T-t}^{i})-\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i}))$
$=\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i})+w(\mathbf{X}_{0}^{i})-\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i}))$
$=\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))-\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}$
(C.10) $\leq\frac{1}{2n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}-\frac{1}{2n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2},$
where the inequality holds from $ab\leq\frac{1}{2}a^{2}+\frac{1}{2}b^{2}$. By a similar argument, we find
$\frac{1}{n}\sum_{i=1}^{n}\langle\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle$
$=\frac{1}{n}\sum_{i=1}^{n}\langle\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\rangle$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}$
(C.11) $\leq\frac{1}{2n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}-\frac{1}{2n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}.$
Substituting (C.10) and (C.11) into (C.9) yields
$\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}+\frac{\lambda}{n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}$
$\leq\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}+\frac{\lambda}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}$
$\quad+\frac{2\lambda}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle.$
By taking expectation with respect to $\{(\mathbf{X}_{0}^{i},\mathbf{X}_{T-t}^{i})\}_{i=1}^{n}$, it follows that for each $h_{t}\in{\mathscr{H}}_{t}$,
${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Big]+\lambda{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Big]$
$\leq{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{t})-w(\mathbf{X}_{0}))^{2}\big]+\lambda{\mathbb{E}}_{\mathbf{X}_{t}}\big[\|\nabla h_{t}(\mathbf{X}_{t})-\nabla h_{t}^{*}(\mathbf{X}_{t})\|_{2}^{2}\big]$
$\quad+2\lambda{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle\Big]$
$\leq\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}$
$\quad+2\lambda{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle\Big],$
where the second inequality follows from Proposition 3.4. This completes the proof of (C.5).
Step 2. It follows from $h_{t}^{*}(\mathbf{X}_{T-t}^{i})={\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]$ that
$\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}$
$=\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i})+{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$=\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}+\frac{1}{n}\sum_{i=1}^{n}({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$\quad+\frac{2}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i})).$
Taking expectation with respect to $\{(\mathbf{X}_{0}^{i},\mathbf{X}_{T-t}^{i})\}_{i=1}^{n}$ implies
(C.12) ${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Big]$
$=\underbrace{{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Big]}_{\text{(i)}}+V_{t}^{2}$
$\quad+2\underbrace{{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))\Big]}_{\text{(ii)}},$
where $V_{t}^{2}$ is defined in Proposition 3.4, and we used the fact that $h_{t}^{*}$ is independent of $\{(\mathbf{X}_{0}^{i},\mathbf{X}_{T-t}^{i})\}_{i=1}^{n}$. For term (i) in (C.12), using Lemma C.7 yields
(C.13) ${\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]\leq{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Big]+8\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}).$
For term (ii) in (C.12), by applying Lemma C.11, we find
(C.14) ${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})(w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}])\Big]\leq 2\bar{B}{\mathfrak{R}}({\mathscr{H}}_{t}).$
Plugging (C.13) and (C.14) into (C.12) obtains (C.6). The inequality (C.7) is a direct conclusion of Lemma C.8.
Step 4. It follows from Lemma C.9 that for each $\xi>0$,
$\lambda{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle\Big]$
@@
$\leq\lambda{\mathbb{E}}\big[(\nabla h_{t}^{*},\nabla h_{t}-\nabla\widehat{h}_{t}^{\lambda})_{L^{2}(p_{T-t})}\big]+\frac{8\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})$
@@
$\leq\frac{30}{\xi}\frac{\lambda d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}+\xi\lambda{\mathbb{E}}\big[\|h_{t}-\widehat{h}_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\big]+\frac{8\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})$
$\leq\frac{30}{\xi}\frac{\lambda d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}+\xi\lambda\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\xi\lambda{\mathbb{E}}\big[\|h_{t}^{*}-\widehat{h}_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\big]+\frac{8\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t}),$
where the first inequality is due to Lemma C.9, the second inequality follows Lemma C.6, and the third inequality comes from the triangular inequality. Finally, setting $\xi=\frac{1}{2\lambda}$ yields (C.8).
Finally, combining (C.5), (C.6), (C.7), and (C.8) completes the proof. ∎
C.3  Auxilary lemmas for the oracle inequality
Lemma C.6.
Suppose Assumptions 1 and 2 hold. Assume that $|v_{t}(\mathbf{x})|<\infty$ for each $\mathbf{x}\in{\mathbb{R}}^{d}$. Then for each $\xi>0$,
$|(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}|\leq\frac{30}{\xi}\frac{d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}+\xi\|v_{t}\|_{L^{2}(p_{T-t})}^{2}.$
Proof of Lemma C.6.
By a direct calculation, we have
$-(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}$
$=-\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla v_{t}(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$

### v2

@@ D.1  Oracle inequality of variationally stable Doob’s matching
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}$
$\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{2\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
@@
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}$
$\quad+2\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2},$
where the last inequality holds from the triangular inequality. Using the triangular inequality again, we have
(D.1) ${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]+\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$\leq 2{\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]+2\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$\quad+2\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$\leq\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{4\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+4\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{approximation error}}$
$\quad+\underbrace{160\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+16\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{generalization error}}$
$\quad+\underbrace{6\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+6\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}}_{\text{regularization gap}}.$
Combining Proposition 4.2 and Lemma D.6 yields
[…]
Substituting (D.2) into (D.1) completes the proof. ∎
Lemma D.1.
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6). Then we have
${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]$ $\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
@@
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the notation $\lesssim$ hides absolute constants.
Proof of Lemma D.1.
For any $h_{t}\in{\mathscr{H}}_{t}$, we have
${\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$={\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t})+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$\leq{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t})+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda}),$
where the inequality holds from that fact that $\widehat{h}_{t}^{\lambda}$ is the minimizer of $\widehat{{\mathcal{J}}}_{t}^{\lambda}$ over the hypothesis class ${\mathscr{H}}_{t}$. Taking expectation on both sides of the inequality yields
${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]$ $={\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})\big]+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda}),$
where the equality holds from ${\mathbb{E}}[\widehat{{\mathcal{J}}}_{t}(h_{t})]={\mathcal{J}}_{t}(h_{t})$ for each fixed $h_{t}$, and the inequality is due to $h_{t}\in{\mathscr{H}}_{t}$. By taking infimum on both sides of the inequality with respect to $h_{t}\in{\mathscr{H}}_{t}$, we have
(D.3) ${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]\leq\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]}_{\text{generalization error}}+\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\Big\}}_{\text{approximation error}}.$
The rest of the proof is divided into three steps.
Step 1. Generalization error in (D.3).
For the generalization error in (D.3), we have the following decomposition:
(D.4) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]$
$=\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Bigg]}_{\text{(G1)}}$